OpenStudy (anonymous):

michelle divides her 8 3/4 boxes of cookies between her self and 2 people how many boxes does each person have michelle divides her 8 3/4 boxes of cookies between her self and 2 people how many boxes does each person have @Mathematics

OpenStudy (anonymous):

can u show me how u got it cause my teacher will tell me that

OpenStudy (anonymous):

between her self and 2 people that's 3 people

OpenStudy (anonymous):

\[35/4\div3=35/12\]

OpenStudy (anonymous):

oh i m really so sorry YZK u r ryt

OpenStudy (anonymous):

meek plz recheck it !!

OpenStudy (anonymous):

why is it wrong

OpenStudy (anonymous):

35/4÷3=35/12 this one is correct

OpenStudy (phi):

If you have time, watch this short video, and the one following http://www.khanacademy.org/video/dividing-fractions-example?playlist=Developmental+Math